OpenStudy (anonymous):

Find the coordinates of the points of intersection for the graphs of x^2+2y^2=20 and y=x-2.

jimthompson5910 (jim_thompson5910):

Since y=x-2, we can replace every 'y' in the first equation with x-2. This is because the two are equivalent or the same. So we get this... \[\Large x^2+2y^2=20\] \[\Large x^2+2(x-2)^2=20\] \[\Large x^2+2(x^2-4x+4)=20\] \[\Large x^2+2x^2-8x+8=20\] I'll let you finish